# Agent-as-Judge for Factual Summarization of Long Narratives

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated near-human performance in summarization tasks based on traditional metrics such as ROUGE and BERTScore. However, these metrics do not adequately capture critical aspects of summarization quality, such as factual accuracy, particularly for long narratives (>100K tokens). Recent advances, such as LLM-as-a-Judge, address the limitations of metrics based on lexical similarity but still exhibit factual inconsistencies, especially in understanding character relationships and states. In this work, we introduce NarrativeFactScore, a novel "Agent-as-a-Judge" framework for evaluating and refining summaries. By leveraging a Character Knowledge Graph (CKG) extracted from input and generated summaries, NarrativeFactScore assesses the factual consistency and provides actionable guidance for refinement, such as identifying missing or erroneous facts. We demonstrate the effectiveness of NarrativeFactScore through a detailed workflow illustration and extensive validation on widely adopted benchmarks, achieving superior performance compared to competitive methods. Our results highlight the potential of agent-driven evaluation systems to improve the factual reliability of LLM-generated summaries. [^1] https://github.com/YeonseokJeong/NarrativeFactScore
@@ 1 Introduction
Figure 1: Comparison of factuality evaluation by LLM and Agent Judge with NarrativeFactScore. Given scenes from The Lord of the Rings, the summary incorrectly claims "Sauron is pursuing Gandalf." The LLM Judge assigns 100% factuality score, while our Agent Judge correctly identifies this error through analyzing atomic facts about characters, assigning 75% NarrativeFactScore, with specific feedback.
@@
In our target scenario of summarizing long narratives (> 100K tokens), metrics such as BooookScore Chang et al. (2024) can measure coherence, but evaluating factuality has remained challenging, as it requires not only comparing summaries of, but also understanding complex webs of facts and evolving character relationships across lengthy narratives. Judging the factuality of such long narratives has therefore inevitably relied on costly human evaluations Kim et al. (2024); Subbiah et al. (2024).
@@
More recently, LLM-as-a-Judge metrics Min et al. (2023); Bishop et al. (2024) have reduced human overheads, leveraging LLM to assess the factuality between the input story and its summary, offering a more cost-effective alternative. These metrics split the summary into smaller units, retrieve similar scene from the input story, and quantify factuality by LLM.
@@
However, directly using LLM to evaluate factuality has two limitations. First, as demonstrated by Kim et al. (2024), the LLM judge fails to accurately assess factuality in narratives that require indirect reasoning, such as understanding character relationships or states. For example, in Figure 1, although Sauron is pursuing Frodo in order to obtain the Ring in The Lord of the Rings, the LLM judge inaccurately evaluates the factuality of a summary which incorrectly states that "Sauron is pursuing Gandalf". This limitation arises because an LLM judge cannot consistently reason about character relationships, so we need a consistent CKG.
Second, the LLM judge measures only factuality scores without explanation, making its evaluations less reliable. In contrast, traditional similarity-based metrics such as ROUGE Lin (2004) and BERTScore Zhang et al. (2019) are reliable because their calculation process is interpretable. Therefore, the calculation process of the factuality should also be qualitative and actionable to ensure reliability. These limitations highlight a need for agents that can actively construct and verify their understanding of narratives, rather than resorting to passive comparisons of text.
@@
To address these limitations, we propose an Agent-as-a-Judge Zhuge et al. (2024) framework for evaluating and refining summaries, which operationalizes active narrative understanding with a novel NarrativeFactScore metric. The key driver of our metric is a consistent Character Knowledge Graph (CKG). Our CKG achieves consistency through two key mechanisms: first, we construct a names graph that consolidates character aliases and variations across scenes, and second, inspired by Wang et al. (2023), we perform multiple rounds of relationship extraction and select relationships that appear frequently across scenes as edges. This construction process ensures that only well-supported character relationships are retained. By leveraging this consistent relationship graph when evaluating the factuality, we can accurately assess even complex narrative facts that require understanding intricate character dynamics.
@@
To address the second limitation, NarrativeFactScore evaluates summaries using a combination of retrieval-based verification with explicit feedback. For each statement in the summary, we retrieve relevant scenes and character relationships from our CKG to calculate a factuality score. Since our metric operates autonomously, it is more cost-effective and faster than Human-as-a-Judge. In addition, it offers feedback for low scores, which makes it more reliable than LLM-as-a-Judge metrics. Recognizing the causes of low scores also contributes to generating more accurate summaries through agent-based refinement.
@@
Using NarrativeFactScore provides two key advantages for long narrative summarization. First, it offers a labor-efficient and fast metric that also approximates human evaluation when evaluating the factuality of summaries. Our metric demonstrates a statistically strong correlation with human evaluation, and a test for differences between human evaluation and our metric yielded a p-value of 0.00003. Second, since it provides feedback on factually incorrect parts, agent-based refinement can improve summarization performance. We show that agent-based refinement improves factuality (+14.03), ROUGE (+2.05), and BERTScore (+0.13) on MovieSum Saxena and Keller (2024a), a movie script summarization dataset, and also improves factuality (+12.26), ROUGE (+2.47), and BERTScore (+0.21) on MENSA Saxena and Keller (2024b), a movie scene saliency dataset.
@@ 2.2 Character Knowledge Graph (CKG)
Since characters are integral to narrative Gurung and Lapata (2024), prior work has aimed to construct as graph to easily utilize them. In narrative texts, CKG shows the unidirectional relationship between a subject and an object character. This process is similar to creating a triple (subject-predicate-object) list in knowledge graph construction Chen et al. (2020). Andrus et al. (2022) utilized the OpenIE system Angeli et al. (2015) for story completion and question answering tasks, integrating it with GPT-3 Brown et al. (2020) to enhance its effectiveness. Alternatively, a recent method Zhao et al. (2024) that assembles CKG directly using LLMs offers a more robust approach, as it better captures the nuanced and complex relationships. Our distinction lies not only in constructing CKGs but also in utilizing them to measure and enhance factuality.
@@ 2.3 Summarization Metrics for Evaluating Factuality
In recent research, efforts have been made to evaluate factuality of long documents. LongDocFACTScore Bishop et al. (2024) improves this process by calculating BARTScore Yuan et al. (2021) only on the semantically similar portions of the source text for each summary sentence, making it an effective method for handling long documents. FActScore Min et al. (2023) further enhances factuality evaluation by decomposing text into atomic facts and verifying each with LLM using information retrieved from the knowledge source. Unlikely these metrics, our metric focuses on character relationships to accurately evaluate factuality and provides actionable feedback to refine factually incorrect parts.
@@ 3 Proposed Method
In this section, we elaborate on NarrativeFactScore for evaluating factuality of long narrative summarization.
Figure 2 illustrates three phases of our framework, which will be detailed in later sections, respectively. (1) Constructing a consistent CKG (Section 3.1), (2) Calculating new factuality metrics using the CKG for accurate and interpretable evaluation (Section 3.2), (3) Refining summaries based on agent feedback to improve factuality (Section 3.3).
@@ 3.1 CKG Extraction
We construct a consistent CKG, to overcome the inconsistencies of CKG reported in Kim et al. (2024); Zhao et al. (2024), losing information Liu et al. (2024) in long narratives and failing to reason over many implicit relationships at once. To address these issues, we perform reasoning multiple times Wang et al. (2023) for each scene and select frequent relationships to improve consistency and accuracy.
@@
  <predicate>Dark Lord
@@
Figure 3: (a) Part of a knowledge graph generated from The Lord of the Rings, with three named entities. ‘Frodo/Frodo Baggins’ is a single entity with two names. (b) The same graph in linearized form.
First, to maintain consistency in character identification, we construct a names graph, consolidating aliases or variations in names in scenes. For example, in The Lord of the Rings, ‘Frodo’ and ‘Frodo Baggins’ are recognized as the same character. As illustrated in Figure 3(a), each variation in the name is a node, with undirected edges connecting the nodes that refer to the same character. This step ensures an accurate capture of relationships, even when names vary across scenes. The knowledge graph is initialized using names from the names graph.
Second, to preserve the consistency of relationships, the extracted triples are processed to select frequent as the final edges. Only triples with named entities as subjects and objects are used; if an object is missing, a self-loop is added to represent the state of character. For triples with the same subject and object, frequent predicates are added as directed edges in the knowledge graph. Since relationships can involve multiple predicates, all predicates that exceed a set threshold are included. Adjusting the threshold allows for control over the graph: a higher threshold ensures greater consistency, while a lower threshold increases diversity. The process of deciding edges is repeated to construct a CKG that can effectively evaluate the factuality of summaries.
@@ 3.2 NarrativeFactScore Calculation
Using the retrieved information, each atomic fact $a_{k}$ is evaluated to determine its factuality and to obtain feedback supporting the evaluation. This evaluation is carried out using the prompt detailed in Section D.4, which produces 1 if the fact is accurate. If the fact is determined to be inaccurate, the prompt also provides feedback $f_{i}$ on how to correct it. Finally, the NarrativeFactScore is calculated as the proportion of atomic facts that are found to be factual, defined by the following equation:
[…]
where $\mathbbm{1}$ is the indicator function, yielding 1 if the atomic fact $a_{i}$ is factual and 0 otherwise.
@@ 3.3 Agent-based Fact Refinement
The new metric leveraging consistent CKG enables the agent to guide refinement by using feedback from the evaluation. This process involves three key inputs: original narrative to provide global context, the initial summary that requires modification, and the feedback detailing the inaccuracies and reasons for those errors. Using these inputs and using the prompt described in Section D.5, the LLM generates an improved summary $y$ that corrects the factual inaccuracies identified earlier.
Motivated by Madaan et al. (2024), the improved summary can be further evaluated as outlined in Section 3.2. This allows the agent-based refinement to be iterative, where each iteration further refines the summary by addressing any remaining inaccuracies, ultimately enhancing the overall factuality.
4 Demonstration Scenarios
This section describes our demonstration scenario, inspired by the real-life workflow of a company collecting and analyzing scripts. In the real world, narratives such as dramas and movies often exceed 100K tokens, which makes manual analysis difficult. Automating the evaluation of summaries is essential not only for summarization itself but also for tasks such as similar content recommendation and narrative-based QA. Summaries are also crucial for deciding whether to watch dramas or movies, which makes long narrative summarization a foundational task for companies in the narrative industry. [^2] The narrative industry broadly refers to the sector that creates, distributes, and analyzes various forms of narratives, such as films, television shows, books, video games, and other media that tell stories. This includes businesses involved in producing, editing and consuming these forms of content, focusing on storytelling in both traditional and digital media. However, people are sensitive to factuality in narrative summarization, yet there is no automated metric to evaluate it. To address this gap, we propose a demo that simplifies access to factuality assessment.
Figure 4 shows screenshots of the system, aligned with the three phases of our framework in Figure 2.333huggingface.co/spaces/yeonseokjeong/NarrativeFactScore Using the example Black Panther, users can view the original narrative after selecting a dataset, data type, and name. Clicking “Generate Knowledge Graph" generates and visualizes the CKG (Section 3.1). The “Generate Initial Summary" and “Calculate Factuality Score" buttons create a initial summary and evaluate its factuality using the CKG (Section 3.2). Finally, “Refine Summary" improves the summary based on feedback, enhancing factuality (Section 3.3).
@@ 5.2 Correlation with Human Factuality Scores
Table 1: Spearman and Kendall’s tau correlation coefficients between different metrics and human factuality assessments on STORYSUMM Subbiah et al. (2024) and FABLES Kim et al. (2024).

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated near-human performance in summarization tasks based on traditional metrics such as ROUGE and BERTScore. However, these metrics do not adequately capture critical aspects of summarization quality, such as factual accuracy, particularly for long narratives (>100K tokens). Recent advances, such as LLM-as-a-Judge, address the limitations of metrics based on lexical similarity but still exhibit factual inconsistencies, especially in understanding character relationships and states. In this work, we introduce NarrativeFactScore (NFS), the first “Agent-as-a-Judge” framework that evaluates and refines factuality in narrative summarization. By leveraging a Character Knowledge Graph (CKG) extracted from input narrative, NarrativeFactScore evaluates the factuality and provides actionable guidance for refinement, such as identifying missing or erroneous facts. Our experimental results demonstrate that constructing the CKG enables reasoning with 1/3 of the factuality computation used in prior approach, and achieve three times higher correlation with human judgments. Furthermore, refinement with actionable guidance improves the quality of the summary. [^1] https://github.com/YeonseokJeong/NarrativeFactScore
@@ 1 Introduction
Figure 1: Comparison of factuality evaluation by LLM and Agent Judge with NarrativeFactScore. Given scenes from The Lord of the Rings, the summary incorrectly claims “Sauron is pursuing Gandalf.” The LLM Judge assigns 100% factuality score, while our Agent Judge correctly identifies this error through analyzing atomic facts about characters, assigning 75% NarrativeFactScore, with specific feedback.
@@
In our target scenario of summarizing long narratives (> 100K tokens), metrics such as BooookScore Chang et al. (2024) can measure coherence, but evaluating factuality has remained challenging Subbiah et al. (2024). This is because it requires comparing summaries not only against complex facts but also against the evolving relationships among characters in long narratives. Thus, judging the factuality of such long narratives has therefore inevitably relied on costly human evaluations Kim et al. (2024).
@@
More recently, LLM-as-a-Judge metrics Min et al. (2023); Bishop et al. (2024) have leveraged LLM to assess the factuality, offering a more cost-effective alternative to human annotations. If applied to narrative summarization, these metrics split the summary into smaller units, retrieve similar scenes from the input story, and quantify factuality by LLM.
@@
However, directly using LLM to evaluate factuality has two limitations. First, as demonstrated by Kim et al. (2024), the LLM judge often fails to accurately assess factuality in narratives that require indirect reasoning, such as understanding character relationships or states. For example, in Figure 1, although Sauron is pursuing Frodo in order to obtain the Ring in The Lord of the Rings, the LLM judge inaccurately evaluates the factuality of a summary which incorrectly reports that “Sauron is pursuing Gandalf”. This limitation stems from the inability of the LLM judge to consistently track and reason about character relationships, highlighting the need to maintain structured and consistent CKG.
Second, the LLM judge evaluates factuality as a score with no interpretability, whereas ROUGE Lin (2004) and BERTScore Zhang et al. (2019) enable to interpret why a given score was assigned and further facilitates the generation of actionable feedback. Desirably, evaluation metrics for summary can provide feedback, when the score is low, to explain why it is incorrect and suggest how to improve.
@@
We propose an Agent-as-a-Judge Zhuge et al. (2024) framework, using interpretable evaluation of summaries with a novel NarrativeFactScore, based on which we can refine and improve summary quality. CKG achieves consistency by constructing a names graph that consolidates character aliases and variations across scenes and by performing multiple rounds of relationship extraction, selecting relationships that frequently appear across scenes as edges, inspired by Wang et al. (2023). This construction process ensures that only well-supported character relationships are retained. By leveraging this consistent relationship graph when evaluating the factuality, we can accurately assess even complex narrative facts that require understanding intricate character dynamics.
@@
To improve the interpretability of the metric, NarrativeFactScore also provides feedback for interpretation and refinement when the summary is incorrect. For each statement in the summary, our metric retrieves relevant scenes and character relationships from our CKG to calculate a factuality score. Based on the retrieved evidence, ours evaluates each statement and generates feedback identifying discrepancies between claims and supporting evidence. Since our metric operates autonomously, it is more cost-effective and faster than Human-as-a-Judge. In addition, it offers feedback for low scores, which makes it more reliable than LLM-as-a-Judge metrics. Recognizing the causes of low scores also contributes to generating more accurate summaries through agent-based refinement.
@@
Using NarrativeFactScore provides two key advantages for long narrative summarization. First, it offers a labor-efficient and fast metric that also approximates human evaluation when evaluating the factuality of summaries. Our metric demonstrates a statistically strong correlation with human evaluation, and a test for differences between human evaluation and our metric yielded statistically significant results, with the p-value falling below 0.05. Second, since it provides feedback on factually incorrect parts, agent-based refinement can improve summarization performance. We show that agent-based refinement improves factuality (+14.03), ROUGE (+2.05), and BERTScore (+0.13) on MovieSum Saxena and Keller (2024a), a movie script summarization dataset, and also improves factuality (+12.26), ROUGE (+2.47), and BERTScore (+0.21) on MENSA Saxena and Keller (2024b), a movie scene saliency dataset.
@@ 2.2 Character Knowledge Graph (CKG)
Since characters are integral to narrative Gurung and Lapata (2024), prior work has aimed to construct a graph to easily utilize them. In narrative texts, CKG shows the unidirectional relationship between a subject and an object character. This process is similar to creating a triple (subject-predicate-object) list in knowledge graph construction Chen et al. (2020). Andrus et al. (2022) utilized the OpenIE system Angeli et al. (2015) for story completion and question-answering tasks, integrating it with GPT-3 Brown et al. (2020) to enhance its effectiveness. Alternatively, a recent method Zhao et al. (2024) that assembles CKG directly using LLMs is more robust approach, as it better captures the nuanced and complex relationships. Our distinction lies not only in constructing CKGs but also in utilizing them to measure and enhance factuality.
@@ 2.3 Summarization Metrics for Evaluating Factuality
In recent research, efforts have been made to evaluate factuality of long documents. LongDocFACTScore Bishop et al. (2024) improves this process by calculating BARTScore Yuan et al. (2021) only on the semantically similar portions of the source text for each summary sentence, making it an effective method for handling long documents. FActScore Min et al. (2023) further enhances factuality evaluation by decomposing text into atomic facts and verifying each with LLM using information retrieved from the knowledge source. Unlike these metrics, our metric focuses on character relationships to accurately evaluate factuality and provide actionable feedback to refine factually incorrect parts.
@@ 3 Proposed Method
In this section, we elaborate on NarrativeFactScore for evaluating factuality of long narrative summarization. Figure 2 illustrates three phases of our framework, which will be detailed in Section 3.1, 3.2, and 3.3 respectively.
@@ 3.1 CKG Extraction
We construct a consistent CKG, to overcome the inconsistencies of CKG reported in Kim et al. (2024); Zhao et al. (2024), losing information Liu et al. (2024) in long narratives and failing to reason over many implicit relationships at once. To address these issues, we perform reasoning multiple times Wang et al. (2023) for each scene and select frequent relationships to improve consistency and accuracy. We note this requires only 1/3 of the original cost while improving correlation threefold. (See Appendix 5.4.)
@@
  <predicate>Dark Lord, desire Ring
@@
Figure 3: (a) Part of a knowledge graph generated from The Lord of the Rings, with three named entities. ‘Frodo/Frodo Baggins’ is a single entity with two names. (b) The same graph is in linearized form.
First, to maintain consistency in character identification, we construct a names graph $G_{name}$, consolidating aliases or variations in names in scenes. Our framework processes each scene in turn, extracts all character names, and determines several names refer to same character based on the context using LLM. For example, in The Lord of the Rings, ‘Frodo’ and ‘Frodo Baggins’ are recognized as the same character. As illustrated by ‘Frodo / Frodo Baggins’ in Figure 3(a), each name variation is a node. [^2] In practice, an undirected edges is added between nodes that refer to the same character. This step ensures an accurate capture of relationships, even when names vary across scenes. The CKG is initialized using names from the names graph.
Second, to preserve the consistency of relationships, we sample extracted triples multiple times Wang et al. (2023); Brown et al. (2024) and select frequent ones as the final edges. Let the node set $V$ be the set of all character in the names graph:
$V=\{v\mid v\in G_{\text{name}}\}$ (1)
Only triples with named entities as subjects and objects are used; if an object is missing, a self-loop is added to represent the state of character. We then define the edge set $E$ of our CKG as
$E=\bigl{\{}(s,p,o)\mid s,o\in V,\ \mathrm{freq}(p\mid s,o)\geq\tau\bigr{\}}$ (2)
where $(s,o)$ denotes a character pair, $\mathrm{freq}(p\mid s,o)$ is the frequency of predicate $p$ for $(s,o)$, and $\tau$ is the frequency threshold. [^3] Adjusting the threshold allows for control over the graph: a higher threshold ensures greater consistency, while a lower threshold increases diversity. Finally, the consistent CKG is given by
$G=(V,E).$ (3)
For triples with the same subject and object, frequent predicates capture temporal changes as directed edges. For example, since the early scenes show that ‘Frodo’ fears ‘Sauron’ and the later scenes show that he resists ‘Sauron’, the CKG in Figure 3(a) displays two distinct relationships. The process of deciding edges is repeated to construct a CKG that can effectively evaluate the factuality of summaries.
@@ 3.2 NarrativeFactScore Calculation
Using the retrieved information, each atomic fact $a_{k}$ is evaluated to determine its factuality and to obtain feedback supporting the evaluation. We then define an indicator $I_{k}$ for factuality based on:
$I_{k}=\mathbbm{1}[a_{k}\ \text{is factual given}\ \mathcal{S}_{i},g)]$ (4)
where $\mathbbm{1}$ is the indicator function, yielding 1 if the atomic fact $a_{i}$ is factual and 0 otherwise. This evaluation is carried out using the prompt detailed in Section D.4, which produces 1 if the atomic fact is accurate. If the atomic fact is determined to be inaccurate, then feedback $f_{k}$ is provided on how to correct it. Finally, the NarrativeFactScore is calculated as the proportion of atomic facts that are found to be factual, defined by the following equation:
$\text{{NarrativeFactScore}}=\frac{\sum_{i=1}^{z}I_{k}}{z}$ (5)
@@ 3.3 Agent-based Fact Refinement
The new metric leveraging consistent CKG enables LLM agent to guide refinement by using feedback from the evaluation. This process involves three key inputs: original narrative to provide global context, the initial summary that requires modification, and the feedback detailing the inaccuracies and reasons for those errors. Using a prompt that incorporates these inputs in Section D.5, LLM refines the initial summary by correcting only the factually inaccurate parts provided from feedback, then generates the improved summary $y$.
Motivated by Madaan et al. (2024), the improved summary can be further evaluated as outlined in Section 3.2. This allows the agent-based refinement to be iterative, where each iteration further refines the summary to enhance overall factuality.
@@ 4 Experiments
4.1 Implementation Details
Uniform Language Model Usage
To ensure that performance gain is due to our framework and not the underlying language model, we use only gpt-4o-mini-2024-07-18 in our experiments. This model is applied across all components, including CKG extraction, summarization, fact decomposition, fact check, and agent-based fact refinement. This approach prevents superior LLMs from influencing the results, allowing us to rigorously evaluate the effectiveness of our framework.
Generating Initial Summary
To generate the initial summary $Z$, we adopt hierarchical merging Chang et al. (2024) that ensures the logical structure of the narrative is preserved. The narrative is first divided into chunks $C_{i}$ where each chunk is formed incrementally by adding scenes until a predefined context size444we set predefined context size of a chunk to 1024. is reached. Once this limit is exceeded, a new chunk begins, resulting in a sequence of chunks $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$. Each chunk $C_{i}$ is then independently summarized using the prompt specified in Section D.2, and the resulting chunk summaries are sequentially merged to produce the initial summary $Z$.
Retrieving Relevant Scene and Subgraph
Using the BGE-M3 embedding model Chen et al. (2024), we retrieve information relevant to each atomic fact $a_{k}$. Specifically, we identify the most similar scene $\mathcal{S}_{i}$ from the narrative $\mathcal{N}$ and a subgraph containing the three most relevant triples in the linearized CKG $G$. All retrieval computations were performed on a single NVIDIA RTX 3090 GPU.
@@ 4.2 Evaluation Metrics
We report ROUGE and BERTScore as reference points for lexical and semantic similarity, while emphasizing that these metrics were not designed to capture factual accuracy. Our primary factuality comparisons are instead carried out with dedicated metrics such as FActScore Min et al. (2023) and LongDocFACTScore Bishop et al. (2024).
@@ 4.3 Correlation with Human Factuality Scores
Table 1: Spearman and KENDALL’s tau correlation coefficients between different metrics and human factuality assessments on STORYSUMM and FABLES. Coefficients indicating strong correlation are underlined. [^6] We follow widely adopted interpretations reported in Table 4.
Metrics STORYSUMM FABLES
Spearman KENDALL Spearman KENDALL
(A) NFS 0.43 0.31 0.47 0.33
(B) 0- consistency 0.21 0.14 0.19 0.13
(C) 0- CKG 0.30 0.21 0.25 0.16
Table 2: Ablation results on STORYSUMM and FABLES, showing the impact of using different CKG.
Dataset
To evaluate whether the NarrativeFactScore we proposed correlates effectively with human factuality, we conducted a series of experiments. For this purpose, we used STORYSUMM  Subbiah et al. (2024) and FABLES Kim et al. (2024), which include multiple summaries generated by LLM for each narrative. These summaries were then evaluated by human annotators based on their factual accuracy.
Results
We computed the Spearman Spearman (1961) correlations and KENDALL’s tau KENDALL (1938) correlations for each metric in relation to the human factuality scores, as shown in Table 6. NarrativeFactScore is the only metric that shows a strong correlation with human annotations in all datasets. This correlation is statistically significant, with p-values below 0.05 for all datasets.
Ablation Study
To verify the effectiveness of CKG in evaluating factuality, we conducted an ablation study in Table 2. NarrativeFactScore (A) iteratively reasons about character relationships, selects frequent relationships to construct a consistent CKG, and utilizes it for factuality evaluation. In contrast, (B) generates the CKG by reasoning character relationships in a single step and evaluates factuality accordingly. However, according to Zhao et al. (2024), LLMs tend to generate inaccurate character relationships when reasoning over long narrative in a single step. Lastly, (C) evaluates factuality without utilizing a CKG.
The experimental results show that our metric (A) achieves the highest correlation with human, and indicate the following observations. First, comparing (A) with (C) shows CKG contributes to more accurate factuality evaluation. However, the results of (B) and (C) show that an inaccurate CKG can hinder factuality evaluation rather than improve it. Thus, to effectively assess the factuality of summary, it is necessary to construct a consistent CKG through multiple iterations of reasoning.
4.4 Summarization Performance Evaluation